40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 9 01:29:26 2017, gautam, Update, LSC, DRMI Nosie Budget v3.0 7x
    Reply  Mon Oct 9 11:55:01 2017, Koji, Update, LSC, DRMI Nosie Budget v3.0 
       Reply  Mon Oct 9 22:18:34 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_dark.pdf
          Reply  Tue Oct 10 22:04:06 2017, rana, Update, LSC, AS55Q Dark Noise 
          Reply  Wed Oct 11 14:42:03 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
             Reply  Wed Oct 11 19:31:32 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
                Reply  Thu Oct 12 01:50:11 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_dark.png
                   Reply  Thu Oct 12 12:17:28 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_AS55_40m.pdf
                      Reply  Fri Oct 13 12:26:12 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_DemodNoises.pdf
                         Reply  Mon Oct 16 16:01:04 2017, gautam, Update, LSC, AS55Q Dark Noise DemodBoardwOscope.pdfDemodBoardwOscope_ASD.pdf
                            Reply  Tue Oct 17 19:31:53 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_TFs.pdfAS55_DemodNoises.pdf
    Reply  Tue Nov 7 17:45:05 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 SRCL_MICH_whitenoise_tf.pdfPRCL_MICH_whitenoise_tf.pdfMICH_aux.pdfC1NB_disp_40m_MICH_NB_2017-10-08.pdf
       Reply  Tue Nov 7 22:56:21 2017, gautam, Update, LSC, DRMI locking recovered DRMI_7Nov20178.png
          Reply  Wed Nov 8 00:28:16 2017, gautam, Update, LSC, Laser intensity coupling measurement attempt PSL_RIN.pdf
       Reply  Wed Nov 8 09:37:45 2017, rana, Update, LSC, DRMI Nosie Budget v3.1 
          Reply  Wed Nov 8 09:59:12 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 
             Reply  Wed Nov 15 01:37:07 2017, gautam, Update, LSC, DRMI low freq. nosie improved C1NB_disp_40m_MICH_NB_2017-11-15.pdfOL_A2L_couplings.pdf
                Reply  Thu Nov 16 00:53:26 2017, gautam, Update, LSC, DRMI noise sub-budgets C1NB_a2l_40m_MICH_NB_2017-11-15.pdfC1NB_aux_40m_MICH_NB_2017-11-15.pdf
Message ID: 13382     Entry time: Mon Oct 16 16:01:04 2017     In reply to: 13380     Reply to this: 13384
Author: gautam 
Type: Update 
Category: LSC 
Subject: AS55Q Dark Noise 

Koji suggested looking at the output of the AS55 demod board on a fast oscilloscope to look for differences in the two channel outputs (if there is some high-frequency oscillations, for example, we could miss this information in the SR785 spectra). Besides, I was only looking at spectra out to a few kHz on the SR785. I grabbed this data with a 300MHz BW Tektronix oscilloscope (battery mode) today. Input impedance of both channels were set to 1Mohm, and the measurement was made with the RFPD input terminated, output of the daughter board is what is measured. The vertical scaling of the channels was set to the minimum allowed, 1mV/div.

Attachment #1 shows that there is indeed a visible difference between the two channels - the (noisier) I channel has a much larger DC offset of ~5mV compared to the Q channel (I tried switching channels on the O'scope and the larger DC offset remained on the I channel, so seems real). There is also some kind of oscillation going on in the I channel, although the frequency is pretty low, with the peaks spaced ~50us apart. Indeed, in the ASD of the acquired data, the excess power in the I channel at 20kHz and higher harmonics are evident (see Attachment #2). Anyway all of this points to something being anomalous on the daughter board I channel signal path - I will pull it out and monitor the outputs at various points along the signal path with the fast scope to see if I can narrow down what's going on where.

Quote:

Both channels should be identical - I don't understand why the I channels are noisier than their Q counterparts. This is almost certainly a problem on the daughter board, as the orange traces are pretty much identical for both channels.

 

Attachment 1: DemodBoardwOscope.pdf  265 kB  | Hide | Hide all
DemodBoardwOscope.pdf
Attachment 2: DemodBoardwOscope_ASD.pdf  178 kB  Uploaded Mon Oct 16 23:43:29 2017  | Hide | Hide all
DemodBoardwOscope_ASD.pdf
ELOG V3.1.3-